OpenStudy (anonymous):

Owen scored ninety-five points on a one hundred-point test and twenty-nine on a thirty-point quiz. On which did he get a larger fraction of the total points correct? Show your work.

OpenStudy (anonymous):

29*(100/30) = 96.66666

OpenStudy (anonymous):

95/100=0.95 29/30=0.966666 So the 29/30 is higher